doglover12:

Now you’ll complete the temperature part of the problem again, but this time using decimal numbers. Start by converting the temperatures in the problem statement from fractions to decimal numbers.

Laylalyssa:

Maybe provide the entire problem I dont understand what this is asking me to do..

doglover12:

ok give me a sec

doglover12:

63 1/4 - 2 1/5 for the next. Write an expression to represent the water temperature after the submarine’s final dive. Then simplify the expression to find the temperature.

supie:

@Laylalyssa

supie:

He's offline actually so don't even bother.

doglover12:

jhonyy9 r u going to replie

jhonyy9:

ok so like a first idea how you convert the 63 1/4 and the 2 1/5 in fractions without integer ? do you know the way ?

doglover12:

no

jhonyy9:

ok here is an example 2 1/3 = (2*3 +1)/3 = 7/3